# Logical Phase Transitions: Understanding Collapse in LLM Logical Reasoning

### v1

@@ 1 Introduction
Symbolic logical reasoning refers to the ability to draw correct conclusions by applying explicit logical rules to structured premises (formal-logic-2003introduction). It is a foundational aspect of human cognition and a core capability expected of large language models (LLMs) (AGI-2020asymptotically; Attention-2017attention). It supports a wide spectrum of tasks, including commonsense inference (Commonsense-reasoning-2024-candle), mathematical proof (Mathematical-reasoning-2023mathcoder; Geometric-reasoning-2024deep), and philosophical analysis (philosophical-thinking-2013critical).
Recent studies (Symbolic-COT-2024faithful; Baseline-Aristotle2024xu; LogicLM-2023) indicate that current LLMs perform well on logical reasoning tasks in simple scenarios, yet experience substantial degradation as reasoning becomes more challenging. While this gap is widely observed, how logical depth shapes reasoning ability in LLMs remains poorly understood, as there is still no clear characterization of reasoning behavior across increasing levels of logical complexity.
@@ 2 Related Work
LLM Logical Reasoning. Prior work spans three main directions: (1) Linear Reasoning (LR) methods such as Naive Prompting and Chain-of-Thought (COT-2022chain); (2) Aggregative Reasoning (AR) approaches (CLOVER-ICLR-2024divide; Cumulative-reasoning-2023-TsinghuIIIS; TOT-2023tree; Determlr-ACL-2024-RenDaGaoling) that combine multiple reasoning trajectories; and (3) Symbolic Reasoning (SR) frameworks (NL2FOL-2023harnessing; Baseline-Aristotle2024xu; Symbolic-COT-2024faithful; LogicAgent-2025-ours) that integrate LLMs (Gpt4-2023gpt; DeepseekR1-2025deepseek; LLM-survey-2023survey) with explicit logic modules (LogicLM-2023). While these methods improve surface performance through guided or modular reasoning, they offer limited insight into how reasoning behavior changes with increasing logical complexity.
Logical Reasoning Capacity Analysis. Recent work analyzes LLM reasoning fidelity, compositionality, and failure modes (Reasoning-survey-2022towards; multi-step-2024exploring). Studies such as InfoQA (InfoQA-ICLR2025-XiuyingChen-MBZUAI), CoT-Valve (CoT-Valve-ACL2025-NUS), and symbolic Monte Carlo supervision (Sym-MC-EMNLP2025-Sheffield) investigate capacity limits, chain-length controllability, and the role of symbolic trajectories, while Apple (illusion-of-thinking-Nips2025-Apple) shows that frontier LMMs suffer sharp reasoning collapses in tasks such as Tower of Hanoi as complexity increases. However, these analyses rely on coarse difficulty proxies and do not provide a principled framework for quantifying logical complexity itself or characterizing collapse behaviors.
@@ 3.1 Logical Complexity Measuring
Building on the data-construction principles (Benchmark-ProverQA-ICLR-2025large), we construct a Neuro-Symbolic Alignment Dataset for Logical Reasoning (NSA-LR) that provides paired natural language (NL) and FOL representations for every sample. All NL propositions, premises, and reasoning steps are translated into explicit predicates, quantifiers, connectives, and multi-step reasoning chains, following rules in Appendix Table 11. And each statement is independently translated by GPT-5 and Qwen3-Max Qwen3-2025qwen3; matching outputs undergo CFG validation, while mismatches are manually adjudicated. A comparison with existing benchmarks, including ProofWriter (Benchmark-ProofWriter-2020proofwriter), ProntoQA (Benchmark-Pronto-2022language), FOLIO (Benchmark-FOLIO2022folio), and ProverQA (Benchmark-ProverQA-ICLR-2025large), is provided in Table 1.
@@ 3.2 Discovery of LPT
Analogous to phase transitions in physics phase-transitions-1937theory, where a system exhibits macroscopic changes once a control variable enters a critical region, we term this phenomenon a Logical Phase Transition (LPT) and formalize it as a collapse in model performance governed by logical complexity. Importantly, the transition is not characterized by a single threshold, but occurs over one or more critical intervals
@@
The discovery of LPTs indicates that direct exposure to high-complexity samples is ineffective, motivating curriculum learning curriculum-learning-2021survey that organizes samples from easier to harder ones. By progressively increasing logical complexity, curriculum learning enables stable traversal of transition regions beyond the LPT regimes.
Experience Pool. Using the critical intervals from the phase-transition analysis, the dataset is stratified into three regimes: Easy samples satisfy $\mathrm{LoCM}(\phi)<\tau^{(1)}_{\min}$; Hard samples satisfy $\mathrm{LoCM}(\phi)>\tau^{(K)}_{\min}$, where $K$ is the number of transition intervals; all remaining samples fall into the Medium regime. This structured Experience Pool provides the basis for complexity-aware curriculum scheduling curriculum-learning-2021survey in the Neuro-Symbolic curriculum tuning.
@@ 3.3 Neuro-Symbolic Curriculum Tuning
Hybrid reasoning in LogicAgent LogicAgent-2025-ours shows that NL and FOL provide complementary strengths: NL contributes semantic grounding while FOL supplies precise symbolic constraints. Their integration consistently yields higher reasoning accuracy than either modality alone, motivating our hybrid-pretraining strategy.
@@ 4 Experiment
4.1 Settings
Symbolic Toolkit. To verify the syntactic validity of FOL forms, we employ the nltk (tool-nltk-2006nltk) library for CFG-based structural checking.
Benchmarks. We evaluate on four established logical reasoning benchmarks: ProntoQA (Benchmark-Pronto-2022language) (5-hop subset), ProofWriter (Benchmark-ProofWriter-2020proofwriter) (depth-5, OWA setting), FOLIO (Benchmark-FOLIO2022folio) (full expert-curated split), and ProverQA (Benchmark-ProverQA-ICLR-2025large) (1500 examples in total, 1–9 reasoning steps). Detailed benchmark descriptions are provided in Appendix B.1.
@@ 4.1 Settings
Baselines. We evaluate six model variants built upon Qwen2.5–7B: (1) the original model (Orig.), (2) a model fine-tuned on ProntoQA data, (3) a model fine-tuned on ProofWriter data, (4) a model fine-tuned on FOLIO data, (5) a model fine-tuned on ProverQA data, and (6) our multi-stage fine-tuned model ($\color[rgb]{1,0.5,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.5,0.0546875}\theta^{*}$}). For each model variant, we compare two prompting baselines: Naive Prompting and Chain-of-Thought (CoT) (COT-2022chain). Detailed descriptions are provided in Appendix B.4.
@@ 4.2 Evaluation Results
Table 2 reports the evaluation results and we can draw two main observations.
@@
(2) Our method on NSA-LR generalizes to heterogeneous reasoning datasets. Our model $\color[rgb]{1,0.5,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.5,0.0546875}\theta^{*}$}, trained via Neuro-Symbolic Curriculum Tuning on the NSA-LR dataset, achieves strong performance on ProntoQA, ProofWriter, FOLIO, and ProverQA, consistently ranking among the top results. In contrast, alternative fine-tuning methods that are optimized on individual datasets exhibit limited gains on their training datasets and often suffer noticeable performance degradation when evaluated on other reasoning datasets, failing to generalize across heterogeneous tasks. These results indicate that curriculum tuning guided by neuro-symbolic alignment captures transferable reasoning structure, enabling performance improvements that extend beyond the training dataset to diverse logical forms and complexity regimes.
@@ A.1 Benchmarks for Logical Reasoning
Existing logical reasoning benchmarks fall broadly into two categories. Synthetic symbolic benchmarks such as PrOntoQA (Benchmark-Pronto-2022language) and ProofWriter (Benchmark-ProofWriter-2020proofwriter) evaluate deduction in controlled symbolic settings using rule-based or template-driven generation. Natural-language logical benchmarks such as FOLIO (Benchmark-FOLIO2022folio) and ProverQA (Benchmark-ProverQA-ICLR-2025large) provide English scenarios paired with verified reasoning chains produced via theorem-proving pipelines. Together, these datasets offer structured environments for assessing the logical validity of model outputs.
@@ A.2 LLM-Based Logical Reasoning
Reasoning Methodologies. Prior work on LLM hu2025sf2t; ye2025mvp logical reasoning can be grouped into three main directions: (1) Linear Reasoning (LR) methods such as Naive Prompting and Chain-of-Thought (COT-2022chain), which generate step-by-step natural-language explanations; (2) Aggregative Reasoning (AR) approaches (CLOVER-ICLR-2024divide; Cumulative-reasoning-2023-TsinghuIIIS; TOT-2023tree; Determlr-ACL-2024-RenDaGaoling) that combine multiple reasoning trajectories through sampling, search, or majority voting; and (3) Symbolic Reasoning (SR) frameworks (NL2FOL-2023harnessing; Baseline-Aristotle2024xu; Symbolic-COT-2024faithful; LogicLM-2023) that integrate LLMs (Gpt4-2023gpt; DeepseekR1-2025deepseek; LLM-survey-2023survey; qwen2.5-2025qwen2) with explicit logic modules or FOL-based reasoning. Together, these methodologies improve performance on a wide range of reasoning benchmarks through guided, multi-path, or logic-augmented reasoning.
Open Challenges. Despite these advances, many methodologies primarily improve prompting, reasoning-time scaffolds, or external reasoning modules, rather than explicitly characterizing or strengthening robustness across varying levels of logical complexity. As a result, most approaches improve accuracy on fixed benchmarks but provide limited insight into how internal reasoning behavior changes along a fine-grained complexity continuum, which is the central focus of our work.
A.3 Logical Reasoning Capacity Analysis
Existing Capacity and Fidelity Analyses. A substantial body of work investigates the reasoning fidelity and compositionality of LLMs GAS3-2025; li2024coupled; song2024autogenic. Prior studies show that models often hallucinate intermediate steps or deviate from valid derivations even when answers are correct (Reasoning-survey-2022towards; multi-step-2024exploring), revealing instability in multi-step reasoning and difficulty maintaining consistent logical chains. More recent capacity-oriented research examines how reasoning behavior changes as tasks grow more complex. InfoQA (InfoQA-ICLR2025-XiuyingChen-MBZUAI) formalizes a single-pass accuracy upper bound for multi-hop QA, demonstrating that performance can collapse once information load exceeds model capacity. CoT-Valve (CoT-Valve-ACL2025-NUS) shows that reasoning chains can be compressed or expanded through latent control, highlighting sensitivity to chain length and internal reasoning load. Symbolic Monte Carlo supervision (Sym-MC-EMNLP2025-Sheffield) leverages symbolic trajectories to stabilize intermediate reasoning. Apple (illusion-of-thinking-Nips2025-Apple) further demonstrates abrupt accuracy collapses in structured puzzles such as Tower of Hanoi as complexity increases.
@@ A.3 Logical Reasoning Capacity Analysis
Appendix B Datasets and Baselines
@@ B.1 Dataset Descriptions
In this section, we describe the logical reasoning datasets and baselines used in our evaluation. The selected datasets span synthetic, tightly controlled environments and expert-curated NL benchmarks, covering diverse logical structures and reasoning demands. Together with the baselines introduced later, this setup enables a systematic and fair assessment of model reasoning behavior across varying levels of deductive complexity.
@@
This benchmark represents a high-quality, expert-curated corpus designed to evaluate natural language reasoning with first-order logic (FOL) annotations. Unlike purely synthetic datasets, FOLIO is authored by human experts, ensuring a high degree of linguistic diversity and naturalness while maintaining strict logical rigor. It requires models to determine the truth value of a conclusion given a set of premises, covering complex logical phenomena such as quantification and negation. The dataset serves as a critical testbed for assessing whether models can handle reasoning tasks that combine rich semantic content with formal logical constraints.
@@
Developed to assess systematic neural logical deduction, ProofWriter features synthetic natural language problems generated over rule-bases under the Open World Assumption (OWA). The dataset requires models to perform multi-step deductive reasoning involving combinations of conjunctions and disjunctions. A key characteristic of ProofWriter is its emphasis on generating proofs that explicitly justify conclusions, thereby allowing for the evaluation of not just the final answer accuracy but also the validity of the underlying reasoning chain.
@@
This synthetic benchmark focuses on analyzing the deductive reasoning capabilities of large language models in a highly controlled symbolic setting. ProntoQA generates examples comprising a set of premises and a step-by-step CoT derivation leading to a final conclusion. By isolating fundamental logical relationships and minimizing linguistic ambiguity, it provides a clean environment for systematically diagnosing the mechanisms of deductive inference and identifying failure modes in multi-hop reasoning.
@@
Built upon the ProverGen framework, ProverQA combines the generative diversity of LLMs with the rigor of automated theorem proving. It contains natural language statements paired with formally verified reasoning chains, designed to test deductive consistency and the alignment between symbolic and linguistic representations. The benchmark is stratified by reasoning depth, offering a scalable environment to evaluate model performance as the number of required reasoning steps increases.

### v2

@@ 1 Introduction
Symbolic logical reasoning refers to the ability to draw correct conclusions by applying explicit logical rules to structured premises (Smith, 2003). It is a foundational aspect of human cognition and a core capability expected of large language models (LLMs) (Cohen et al., 2020; Vaswani et al., 2017). It supports a wide spectrum of tasks, including commonsense inference (Wang et al., 2024; Liu et al., 2026), mathematical proof (Wang et al., 2023; Eisner et al., 2024; An et al., 2025), and philosophical analysis (Paul, 2013; Lin et al., 2026b).
Recent studies (Xu et al., 2024b, a; Pan et al., 2023) indicate that current LLMs perform well on logical reasoning tasks in simple scenarios, yet experience substantial degradation as reasoning becomes more challenging. While this gap is widely observed, how logical depth shapes reasoning ability in LLMs remains poorly understood, as there is still no clear characterization of reasoning behavior across increasing levels of logical complexity.
@@ 2 Related Work
LLM Logical Reasoning. Prior work spans three main directions: (1) Linear Reasoning (LR) methods such as Naive Prompting and Chain-of-Thought (Wei et al., 2022); (2) Aggregative Reasoning (AR) approaches (Ryu et al., 2024; Zhang et al., 2023; Yao et al., 2023; Sun et al., 2024) that combine multiple reasoning trajectories; and (3) Symbolic Reasoning (SR) frameworks (Yang et al., 2023; Xu et al., 2024a, b; Zhang et al., 2025c) that integrate LLMs (Achiam et al., 2023; Guo et al., 2025; Zhao et al., 2023) with explicit logic modules (Pan et al., 2023). While these methods improve surface performance through guided or modular reasoning, they offer limited insight into how reasoning behavior changes with increasing logical complexity.
Logical Reasoning Capacity Analysis. Recent work analyzes LLM reasoning fidelity, compositionality, and failure modes (Huang and Chang, 2022; Zheng et al., 2024; Fang et al., 2025). Studies such as InfoQA (Wan et al., 2025), CoT-Valve (Ma et al., 2025), and symbolic Monte Carlo supervision (Tan et al., 2025) investigate capacity limits, chain-length controllability, and the role of symbolic trajectories, while Apple (Shojaee* et al., 2025) shows that frontier LMMs suffer sharp reasoning collapses in tasks such as Tower of Hanoi as complexity increases. However, their setting focuses on procedural execution in structured puzzles, whereas we study symbolic logical reasoning under explicit propositional/FOL structure; the two works therefore differ in complexity definition, evaluation target, and intervention scope. More broadly, these analyses rely on task-specific or coarse difficulty proxies and do not provide a principled framework for quantifying logical complexity itself or characterizing collapse behaviors.
@@ 3.1 Logical Complexity Measuring
Building on the data-construction principles (Qi et al., 2025), we construct a Neuro-Symbolic Alignment Dataset for Logical Reasoning (NSA-LR) that provides paired natural language (NL) and FOL representations for every sample. All NL propositions, premises, and reasoning steps are translated into explicit predicates, quantifiers, connectives, and multi-step reasoning chains, following rules in Appendix Table 12. And each statement is independently translated by GPT-5 and Qwen3-Max Yang et al. (2025a); matching outputs undergo CFG validation, while mismatches are manually adjudicated. A comparison with existing benchmarks, including ProofWriter (Tafjord et al., 2020), ProntoQA (Saparov and He, 2022), FOLIO (Han et al., 2022), and ProverQA (Qi et al., 2025), is provided in Table 1.
@@ 3.2 Discovery of LPT
Analogous to phase transitions in physics Landau et al. (1937), where a system exhibits macroscopic changes once a control variable enters a critical region, we term this phenomenon a Logical Phase Transition (LPT) and formalize it as a collapse in model performance governed by logical complexity. Importantly, the transition is not characterized by a single threshold, but occurs over one or more critical intervals
@@
The discovery of LPTs indicates that direct exposure to high-complexity samples is ineffective, motivating curriculum learning Wang et al. (2021) that organizes samples from easier to harder ones. By progressively increasing logical complexity, curriculum learning enables stable traversal of transition regions beyond the LPT regimes.
Experience Pool. Using the critical intervals from the phase-transition analysis, the dataset is stratified into three regimes: Easy samples satisfy $\mathrm{LoCM}(\phi)<\tau^{(1)}_{\min}$; Hard samples satisfy $\mathrm{LoCM}(\phi)>\tau^{(K)}_{\min}$, where $K$ is the number of transition intervals; all remaining samples fall into the Medium regime. This structured Experience Pool provides the basis for complexity-aware curriculum scheduling Wang et al. (2021) in the Neuro-Symbolic curriculum tuning.
@@ 3.3 Neuro-Symbolic Curriculum Tuning
Hybrid reasoning in LogicAgent Zhang et al. (2025c) shows that NL and FOL provide complementary strengths: NL contributes semantic grounding while FOL supplies precise symbolic constraints. Their integration consistently yields higher reasoning accuracy than either modality alone, motivating our hybrid-pretraining strategy.
@@ 4 Experiment
Table 3: Comparison with recent training-free reasoning methods on NSA-LR.
Method Low Medium High Overall
Naïve 60.6 49.8 38.6 49.6
+ NSCT 60.9 (+0.3) 50.2 (+0.4) 41.6 (+3.0) 50.9 (+1.3)
CoT 75.5 58.4 39.4 57.7
+ NSCT 84.0 (+8.5) 64.2 (+5.8) 46.8 (+7.4) 65.0 (+7.3)
ToT 82.0 64.8 42.4 63.2
+ NSCT 88.6 (+6.6) 68.3 (+3.5) 48.5 (+6.1) 68.4 (+5.2)
DetermLR 84.0 63.9 43.8 63.9
+ NSCT 87.5 (+3.5) 67.9 (+4.0) 49.2 (+5.4) 68.2 (+4.3)
SymbCoT 83.8 63.7 46.9 64.8
+ NSCT 89.2 (+5.4) 68.5 (+4.8) 50.5 (+3.6) 69.5 (+4.7)
4.1 Settings
Symbolic Toolkit. To verify the syntactic validity of FOL forms, we employ the nltk (Bird, 2006) library for CFG-based structural checking.
Benchmarks. We evaluate on four established logical reasoning benchmarks: ProntoQA (Saparov and He, 2022) (5-hop subset), ProofWriter (Tafjord et al., 2020) (depth-5, OWA setting), FOLIO (Han et al., 2022) (full expert-curated split), and ProverQA (Qi et al., 2025) (1500 examples in total, 1–9 reasoning steps). Detailed benchmark descriptions are provided in Appendix B.1.
Baselines. We evaluate six model variants built upon Qwen2.5–7B: (1) the original model (Orig.), (2) a model fine-tuned on ProntoQA data, (3) a model fine-tuned on ProofWriter data, (4) a model fine-tuned on FOLIO data, (5) a model fine-tuned on ProverQA data, and (6) our multi-stage fine-tuned model ($\color[rgb]{1,0.5,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.5,0.0546875}\theta^{*}$}). For each model variant, we compare two prompting baselines: Naive Prompting and Chain-of-Thought (CoT) (Wei et al., 2022).
@@ 4.2 Evaluation Results
Tables 2 and 3 report the evaluation results, from which we draw three main observations.
@@
(2) Our method on NSA-LR generalizes to heterogeneous reasoning datasets. Our model $\color[rgb]{1,0.5,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.5,0.0546875}\theta^{*}$}, trained via Neuro-Symbolic Curriculum Tuning on the NSA-LR dataset, achieves strong performance on ProntoQA, ProofWriter, FOLIO, and ProverQA, consistently ranking among the top results. In contrast, alternative fine-tuning methods that are optimized on individual datasets exhibit limited gains on their training datasets and often suffer noticeable performance degradation when evaluated on other reasoning datasets, failing to generalize across heterogeneous tasks. These results indicate that curriculum tuning guided by neuro-symbolic alignment captures transferable reasoning structure, enabling performance improvements that extend beyond the training dataset to diverse logical forms and complexity regimes. One exception is FOLIO under CoT prompting, where the improvement is not maintained; we provide further analysis in Appendix C.x.
(3) Our method consistently enhances diverse inference-time reasoning strategies. As shown in Table 3, NSCT improves performance not only under Naive and CoT prompting, but also when combined with recent training-free reasoning methods such as ToT, DetermLR, and SymbCoT. These gains are consistent across all settings, indicating that the effect of NSCT does not depend on a specific prompting format or decoding heuristic. Notably, the improvements are most pronounced in the High-LoCM regime, where logical phase transitions are most severe: under CoT, ToT, and DetermLR, the gains in the High-LoCM band are +7.4, +6.1, and +5.4, respectively, all larger than those in the Low- or Medium-LoCM bands. Overall, these results support the view that NSCT is not merely a narrow fine-tuning heuristic tied to a specific prompting strategy. Instead, it behaves as a transferable training mechanism that consistently improves logical robustness across diverse decoding settings.
@@ 5 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (Numbers 62272184 and 62402189), the China Postdoctoral Science Foundation (Numbers 2024M751012, 2025T180429, and GZC20230894), the Postdoctor Project of Hubei Province (Number 2024HBBHCXB014), the Natural Science Foundation of Hubei Province No.JCZRMS202600758, and Sponsored by CIPS-SMP-Zhipu Large Model Fund (CIPS-SMP20250306), The computation is completed in the HPC Platform of Huazhong University of Science and Technology.
@@ Acknowledgments
The Usage of LLM
In accordance with ACL guidelines, we used large language models solely for writing assistance and language refinement.
@@ A.1 Benchmarks for Logical Reasoning
Existing logical reasoning benchmarks fall broadly into two categories. Synthetic symbolic benchmarks such as PrOntoQA (Saparov and He, 2022) and ProofWriter (Tafjord et al., 2020) evaluate deduction in controlled symbolic settings using rule-based or template-driven generation. Natural-language logical benchmarks such as FOLIO (Han et al., 2022) and ProverQA (Qi et al., 2025) provide English scenarios paired with verified reasoning chains produced via theorem-proving pipelines. Together, these datasets offer structured environments for assessing the logical validity of model outputs.
@@ A.2 LLM-Based Logical Reasoning
Reasoning Methodologies. Prior work on LLM Liu et al. (2025); Hu et al. (2025); Ye et al. (2025); Zeng et al. (2026); Han et al. (2026); Li et al. (2025a) logical reasoning can be grouped into three main directions: (1) Linear Reasoning (LR) methods such as Naive Prompting and Chain-of-Thought (Wei et al., 2022; Chen et al., 2025; Lin et al., 2026a), which generate step-by-step natural-language explanations; (2) Aggregative Reasoning (AR) approaches (Ryu et al., 2024; Zhang et al., 2023; Yao et al., 2023; Sun et al., 2024; Song et al., 2026, 2024a) that combine multiple reasoning trajectories through sampling, search, or majority voting; and (3) Symbolic Reasoning (SR) frameworks (Yang et al., 2023; Xu et al., 2024a, b; Pan et al., 2023) that integrate LLMs (Yang et al., 2025b; Wang et al., 2026b, a; Li et al., 2025b) with explicit logic modules or FOL-based reasoning. Together, these methodologies improve performance on a wide range of reasoning benchmarks through guided, multi-path, or logic-augmented reasoning.
Open Challenges. Despite these advances, many methodologies primarily improve prompting, reasoning-time scaffolds, or external reasoning modules Ling et al. (2026); Chen and Zeng (2025), rather than explicitly characterizing or strengthening robustness across varying levels of logical complexity. As a result, most approaches improve accuracy on fixed benchmarks but provide limited insight into how internal reasoning behavior changes along a fine-grained complexity continuum, which is the central focus of our work.
@@
A.3 Logical Reasoning Capacity Analysis
Existing Capacity and Fidelity Analyses. A substantial body of work investigates the reasoning (Li et al., 2026b, a; Hao et al., 2025, 2026; Zhang et al., 2025a) fidelity and compositionality of LLMs Zhang et al. (2025b); Li et al. (2024); Song et al. (2024b); Chen et al. (2026); Zhang et al. (2026); Zhou et al. (2026). Prior studies show that models often hallucinate intermediate steps or deviate from valid derivations even when answers are correct (Huang and Chang, 2022; Zheng et al., 2024), revealing instability in multi-step reasoning and difficulty maintaining consistent logical chains. More recent capacity-oriented research examines how reasoning behavior changes as tasks grow more complex. InfoQA (Wan et al., 2025) formalizes a single-pass accuracy upper bound for multi-hop QA, demonstrating that performance can collapse once information load exceeds model capacity. CoT-Valve (Ma et al., 2025) shows that reasoning chains can be compressed or expanded through latent control, highlighting sensitivity to chain length and internal reasoning load. Symbolic Monte Carlo supervision (Tan et al., 2025) leverages symbolic trajectories to stabilize intermediate reasoning.
Apple (Shojaee* et al., 2025) further demonstrates abrupt accuracy collapses in structured puzzles such as Tower of Hanoi as complexity increases.
@@ A.3 Logical Reasoning Capacity Analysis
Appendix B Dataset Details
@@ B.1 Dataset Descriptions
We describe the logical reasoning datasets used in our evaluation, which together provide a systematic and fair testbed for assessing model reasoning across diverse logical structures and deductive complexity levels.
@@
FOLIO is an expert-curated benchmark for natural language reasoning with first-order logic (FOL) annotations. Compared with synthetic datasets, it offers greater linguistic diversity and naturalness while preserving strict logical validity. The task requires determining whether a conclusion follows from a set of premises, involving phenomena such as quantification and negation, and thus serves as a key testbed for evaluating the integration of semantic understanding and formal reasoning.
@@
ProofWriter is a synthetic benchmark designed to evaluate systematic neural deduction under the Open World Assumption. It features rule-based natural language problems requiring multi-step reasoning over conjunctions and disjunctions. Its explicit proof supervision enables evaluation of both final-answer correctness and the validity of intermediate deductive steps.
@@
ProntoQA is a synthetic benchmark for diagnosing deductive reasoning in a highly controlled symbolic setting. Each instance contains premises, a step-by-step chain of thought, and a conclusion, allowing precise analysis of multi-hop inference while minimizing linguistic ambiguity.
@@
ProverQA, built on ProverGen, combines LLM-generated linguistic diversity with formally verified theorem-proving chains. It is designed to assess deductive consistency and symbolic–linguistic alignment, with stratified reasoning depths that support scalable evaluation across increasing inference complexity.